• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name yez

    Log entry time 16:05:11 on February 24, 2011

    Entry number 348453

    This entry is a followup to: 348435


    keyword=L-HRS: move two cables of Pion Rejectors again

    I put the cable labeled as 17L~32L back to Roc.3.24:16~31 because it is
    not the cable which gives empty signals (sorry it is my mistake). The
    cable which give issues on prl1.17,18,19,20,21 is actually 1R~16R (the
    map is very ticky and easy to make mistake). I now move this cable from
    Slot24:32~47 to Slot25:32~47.
    I found that there is a label on the rack saying that the ADC card at
    Slot25 has bad channels from 0~31, what explains that we still can not
    see 33,34L/R after David moved them from 0~15 to 16~31. I move this cable
    to Slot25:32~47. Hopefully those channels are good. 

    A copy of this log entry has been emailed to: danez@jlab.org